if TP= 20KL^3 and the producer is purchasing 40 of K and 40 of L , the MRTS IS 1

TRUE OR FALSE

if an industry were to change from monopolistic competitive to perfect competitive, the price of the product would decrease and there would be less product variety

TRUE OR FALSE

if the price is below my average total cost, i should shut down in the long run but not necessarily in the short run

TRUE OR FALSE
Answer #1

1-FALSE

MRTS =MPl/MPk = 60KL^2/20L^3 =3K/L=3

2-TRUE

if an industry were to change from monopolistic competitive to perfect competitive price will decrease because firms don't have market power in perfectly competitive industry and there would be less product variety because of homogeneity of products.

3- FALSE

if the price is below my average total cost, i should exit in the long run.